KOCN HNCO ОН OK

Show curved arrows to indicate the flow of electrons. Show appropriate charges and/or lone electron pairs to initiate electron movement. Show any bonds that must be explicitly drawn when transferring bonding electrons.
Answer #1

[Koeken}-{k®o=c=ije] Кон• Роге - за O® + K + O=C=N-H OK + 0=e=N-H